# 3DGS.zip: A survey on 3D Gaussian Splatting Compression Methods

### v4

@@ 3.2 Structuring and Dimensionality Reduction
While vector quantization focuses on reducing redundancy by encoding attributes into shared clusters, an alternative perspective is to approach compression through structural organization and contextual redundancy exploitation. Instead of encoding attributes in isolation, methods like octrees, hash-grids and self-organizing grids aim to reshape and recombine data spatially. These techniques organize scene elements hierarchically or around representative points, enabling efficient reuse of similar properties across regions. By structuring Gaussians to exploit contextual relationships — such as proximity, color, or shape — compression becomes more about the global arrangement of data, reducing the need to store individual attributes independently. This shift from attribute-focused encoding to structural compression highlights the potential to achieve compact representations through spatial coherence and redundancy management.
3.2.1 Hierarchical and Adaptive Representations
These methods organize 3D Gaussians by exploiting spatial structures, from partitioning with octrees to grouping similar features using anchors, enabling more compact and efficient representations.
@@ 3.2.1 Hierarchical and Adaptive Representations
Octrees are a spatial partitioning technique used in computer graphics to efficiently represent 3D data like point clouds and volumetric models [32]. They have been successfully applied to point cloud compression [41]. Since 3DGS scenes are effectively point clouds with additional attributes, using octrees for 3DGS becomes straightforward. Octrees recursively subdivide 3D space into smaller cubes, allowing efficient memory allocation by focusing on non-empty regions. In addition, coordinates can be referenced relative to sub-cubes leading to a smaller average bit-length. In 3DGS, octrees help allocate memory only to occupied parts of the scene, skipping over empty regions and, reducing memory usage while preserving detail, making them ideal for large-scale scenes.
@@
Anchor-based representations use anchor points as proxies to predict the properties of associated Gaussian kernels. Unlike octrees, which partition space into a hierarchical grid, anchors group Gaussians by associating them with representative points, allowing efficient compression without strict spatial subdivision [29]. Anchors are initialized by voxelizing the 3D scene and are assigned context features, location, scaling, and learnable offsets. By deriving Gaussian attributes from these anchors rather than storing them individually, redundancy is reduced, leading to decreased memory usage. This approach is conceptually related to the codebooks used in vector quantization presented in Section 3.1, where shared representations efficiently reduce storage requirements while retaining high fidelity.
@@
Hash-grid assisted context modeling leverages multi-resolution hash grids, inspired by Instant-NGP [36], to efficiently represent spatial relationships in 3D Gaussian Splatting. Instant-NGP’s multiresolution hash encoding uses a compact hash table to store trainable feature vectors, which are optimized during training to represent complex spatial details with fewer parameters compared to dense grid encodings. This method enables a trade off between memory, performance, and quality by adjusting parameters such as hash table size, feature vector size, and the number of resolutions. Unlike traditional spatial encodings that explicitly manage collisions, the hash-grid approach allows the neural network to implicitly handle these collisions, simplifying implementation and reducing computational overhead. For 3DGS, the features stored in the hash table can be decoded into attributes for splats. Hash grids are particularly effective in allocating memory to regions of high importance, as demonstrated by Instant-NGP, which achieves small, high-quality radiance fields with minimal parameters. This makes them a very effective tool for 3DGS compression.
3.2.2 Dimensionality Reduction Techniques
Dimensional reduction is a key theme for reducing the complexity and improving the efficiency of 3DGS models. The goal of these techniques is to transform 3D data into lower-dimensional forms that are easier to compress and manage, while still retaining essential spatial relationships.
@@ 3.2.2 Dimensionality Reduction Techniques
Z-order curves, or Morton ordering, work by interleaving the bits of the coordinate values from multiple dimensions to create a single one-dimensional index, effectively mapping multidimensional data into a linear sequence while preserving spatial locality. By ordering the Gaussians according to their positions along a Z-order curve, spatial coherence can be exploited to improve the efficiency of run-length or predictive encoding techniques. This makes Z-order curves useful for spatial indexing and creating a coherent order for splats. In the context of sparse point clouds in 3DGS, Z-order curves are used to order splats based on their positions, which is memory efficient as it avoids allocating storage to empty regions. However, this approach has limitations in terms of maintaining neighborhood relationships: high-dimensional neighbors are not always close when mapped linearly, which affects the efficiency of operations that rely on true spatial proximity. More sophisticated methods, like Hilbert curves, can sometimes provide better locality for such datasets[7], offering improved efficiency in spatial indexing by ensuring that spatial neighbors are more likely to remain neighbors in the linear mapping.
@@
Self-Organizing Gaussians are an approach to map high-dimensional Gaussian parameters into 2D grids[35]. The idea is based on the concept of Self-Organizing Maps[22], an unsupervised learning model that projects high-dimensional data onto a lower-dimensional grid, preserving the topological relationships between data points through competitive learning. Similar in motivation to Z-curves, this 2D representation enables the use of standard image compression techniques, exploiting perceptual redundancies and ensuring local smoothness between neighboring splats. While Z-order curves preserve local neighborhoods well, this may break down with the sparse point clouds of 3DGS. Sorting into locally smooth 2D grids ensures that neighbors in both dimensions are similar, providing smoother results. This allows not only the use of run-length or predictive encoding techniques but also enables the use of image coding methods to store the results efficiently. Unlike Z-order curves, which impose a fixed ordering, the Self-Organizing Gaussians representation is optimized for each scene, allowing neighbors in all dimensions to be modeled effectively. By organizing attributes into multiple data layers that share a consistent 2D layout, this technique provides a highly compressible and efficient representation of the original 3D scene.
3.2.3 Advanced Attribute Compression Techniques
Beyond structuring and dimensional reduction, significant gains can be achieved through adaptive compression of Gaussian attributes. These methods focus on reducing the entropy of attributes such as position, rotation, color, and density, while retaining high-fidelity rendering capabilities.
Region Adaptive Hierarchical Transform (RAHT) is an adaptive transformation technique designed to efficiently reduce the entropy of Gaussian attributes in 3D Gaussian Splatting. The primary motivation for RAHT is to capture correlations among neighboring attributes by transforming spatial data hierarchically, which makes them more suitable for entropy coding. RAHT has its roots in point cloud compression, where similar techniques are used to efficiently compress geometric and color information by leveraging local correlations. RAHT works by iteratively partitioning the spatial domain into hierarchical regions, transforming attributes within these regions to maximize redundancy reduction. This process is particularly useful for attributes such as color, density, and other Gaussian properties that often exhibit strong local correlations. By capturing these correlations effectively, RAHT minimizes redundancy, allowing for more compact storage while maintaining the essential spatial and visual information of the scene.
3.2.4 Summary and Connections
The methods discussed form a toolkit to achieve compact and efficient 3D Gaussian Splatting, each with distinct advantages. Z-curves and 2D grid mappings reduce dimensional complexity, but Z-order curves may lose spatial coherence in sparse datasets. Self-Organizing Gaussians overcome this by ensuring smoothness across dimensions, making them suitable for compression using image coding. Octrees adaptively allocate memory only to occupied regions, optimizing resource use for complex scenes. Anchor-based methods group Gaussians around representative points, reducing redundancy efficiently, similar to vector quantization. Hash-grids add adaptive, multiresolution indexing, balancing memory efficiency and rendering quality.
Advanced attribute compression, such as RAHT, captures correlations through hierarchical transformations, effectively minimizing redundancy. RAHT’s origins in point cloud compression help retain essential scene information by exploiting local correlations. Together, these methods enable efficient, high-quality 3D Gaussian Splatting, suitable for real-time applications in constrained environments. Spatial structuring, dimensional reduction, and attribute compression form a comprehensive approach to 3D scene representation.
@@ 3.4 Compaction
By assessing criteria such as gradients, pixel coverage, and saliency maps, ADC intelligently determines whether Gaussians should be cloned, split, or removed. This ensures that additional Gaussians are allocated where they are most needed, such as in high-frequency regions, while redundant or less impactful Gaussians are pruned. Figure 4 exemplifies the Gaussian Pruning process. Based on above mentioned criteria, Gaussians are selected and removed from the scene.
Figure 4: Gaussian pruning process.
@@ 4 Efficient Strategies for Compression and Compaction in 3D Gaussian Splatting
As 3D Gaussian Splatting (3DGS) evolves as a prominent method for real-time scene rendering, its increasing adoption is challenged by substantial storage and computational requirements. This section addresses two essential optimization strategies: Compression (Sec. 4.1) and Compaction (Sec. 4.2). Compression reduces memory usage by employing methods such as vector quantization, which clusters similar Gaussian attributes to reduce redundancy, and structured representations that organize Gaussians into more compact forms like grids or anchor points. Compaction, on the other hand, focuses on optimizing the number and distribution of Gaussians, ensuring that only the most essential elements are retained while reducing unnecessary data. Figure 6 illustrates how Compression – attribute compression and structured representations – and Compaction intersect. Compaction here includes Gaussian pruning and densification approaches. Together, these strategies enhance the efficiency of 3DGS, making it more practical for a wide range of applications and devices.
@@ 4.1.1 Efficient Representation of Gaussian Attributes
Based on the assumption that many Gaussians share similar attributes, they can be quantized using Vector Quantization (VQ). Most commonly quantized attributes are then stored in codebooks based on K-means as in [37, 38, 39, 23, 11]. In LightGaussian [11], the authors employ VQ on Spherical Harmonics in combination with a significance score to omit VQ on Spherical Harmonics with higher significance. Similarly, [38] introduces a sensitivity-parameter which describes the sensitivity of the reconstruction quality to changes of the Gaussian attributes. This sensitivity measure is then used for sensitivity-aware VQ of Gaussian attributes (e.g.  Spherical Harmonics, shape). In [47], the authors categorize SH coefficients in degrees greater than 0 as unimportant and thus compressable through VQ. Furthermore for key attributes (opacity, scales, Euler angles, and 0-degree SH coefficients), Xie et al. reduce the entropy using RAHT before quantization. RDO-Gaussian [43] use entropy-constrained VQ with codebooks to quantize covariance and color parameters for a more compact representation. HAC [8] uses an anchor structure with associated Gaussians and introduces and Adaptive Quantization Module (AQM) designed to dynamically select quantization steps to facilitate entropy coding of the anchor attributes. The researchers in EAGLES [14] employ a methodology for quantizing attributes such as rotation, view-dependent color, and opacity by leveraging a latent vector associated with each attribute. This latent vector is integrated with a multilayer perceptron (MLP) decoder, which facilitates the decoding of latent representations into attribute values. To ensure differentiability during the training process, an additional latent approximation is preserved. A Straight-Through Estimator (STE) is then utilized to round the latent approximation before the propagation of gradients.
As described in Section 2, Gaussian splats are characterized by 59 attributes, with the majority pertaining to color representation. Of these, the Spherical Harmonics (SH) coefficients encompass 48 attributes across three bands of SH. Notably, 45 out of the 48 SH coefficients are responsible for depicting view-dependent colors. To reduce storage, attribute pruning SH coefficients is a common strategy. In light of this, [39] introduces an adaptive adjustment mechanism for SH. The authors advocate for the calculation of an average transmittance per pixel and per view. In instances where there is minimal variation per view, this allows for the reduction of SH coefficients to lower bands or their complete removal, thereby significantly diminishing the memory footprint associated with each Gaussian splat.

### v5

@@ 3.2 Structuring and Dimensionality Reduction
While vector quantization focuses on reducing redundancy by encoding attributes into shared clusters, an alternative perspective integrates structural organization, spatial redundancy exploitation, and dimensional reduction. Instead of encoding attributes in isolation, methods like octrees, hash-grids, and self-organizing grids reshape and recombine data spatially. These techniques organize scene elements hierarchically or around representative points, enabling efficient reuse of similar properties across regions. By structuring Gaussians to exploit contextual relationships—such as proximity, color, or shape—compression becomes more about the global arrangement of data, reducing the need to store individual attributes independently. This shift from attribute-focused encoding to structural compression highlights the potential for compact representations through spatial coherence, redundancy management, and transformations into lower-dimensional forms that preserve essential spatial relationships. In this section, we look at different methods that provide either a re-structuring of the data, a dimensionality reduction or factorization, as well as combinations thereof.
3.2.1 Octrees
@@
Octrees are a spatial partitioning technique used in computer graphics to efficiently represent 3D data like point clouds and volumetric models [37]. They have been successfully applied to point cloud compression [47]. Furthermore, octrees have demonstrated their effectiveness in speeding up Neural Radiance Field (NeRF) rendering through hierarchical volumetric representation, enabling real-time performance [56]. Since 3DGS scenes are effectively point clouds with additional attributes, using octrees for 3DGS becomes straightforward. Octrees recursively subdivide 3D space into smaller cubes (see Figure 4), allowing efficient memory allocation by focusing on non-empty regions. In addition, coordinates can be referenced relative to sub-cubes leading to a smaller average bit-length. In 3DGS, octrees help allocate memory only to occupied parts of the scene, skipping over empty regions and, reducing memory usage while preserving detail, making them ideal for large-scale scenes.
Figure 4: Example of a point cloud with Octree partitioning.
3.2.2 Anchor-based Representations
@@
Anchor points can be used as proxies to predict the properties of associated Gaussian kernels. Unlike octrees, which partition space into a hierarchical grid, anchors group Gaussians by associating them with representative points, allowing efficient compression without strict spatial subdivision [34]. Anchors are initialized by voxelizing the 3D scene and are assigned position, context features, scaling, and learnable offsets. By deriving Gaussian attributes from these anchors rather than storing them individually, redundancy is reduced, leading to decreased memory usage (Fig. 5). This approach is conceptually related to the codebooks used in vector quantization presented in Section 3.1, where shared representations efficiently reduce storage requirements while retaining high fidelity.
Figure 5: Anchor-based structure. Left: A voxalized scene with SfM inistialized points. Right: The center of each voxel becomes an anchor and is associated with position, feature, scaling and offset. From each anchor neural Gaussians are spawned.
3.2.3 Multi-Resolution Hash Grids
@@
Hash-grid assisted context modeling leverages multi-resolution hash grids, inspired by Instant-NGP [41], to efficiently represent spatial relationships in 3D Gaussian Splatting. Instant-NGP’s multiresolution hash encoding uses a compact hash table to store trainable feature vectors, which are optimized during training to represent complex spatial details with fewer parameters compared to dense grid encodings. This method enables a trade off between memory, performance, and quality by adjusting parameters such as hash table size, feature vector size, and the number of hash levels. Unlike traditional spatial encodings that explicitly manage collisions, the hash-grid approach uses an MLP to implicitly handle these collisions. This makes this a hybrid explicit/implicit method, and introduces a small computational overhead for the feature decoding. For application in 3DGS, the features stored in the hash table are decoded into attributes for individual primitives. Hash grids are particularly effective in allocating memory to regions of high importance, as demonstrated by Instant-NGP, which achieves small, high-quality radiance fields with minimal parameters. This makes them a very effective tool for 3DGS compression.
3.2.4 Z-order Curves
@@
Z-order curves, or Morton ordering, work by interleaving the bits of the coordinate values from multiple dimensions to create a single one-dimensional index, effectively mapping multidimensional data into a linear sequence while preserving spatial locality. Figure 6 exemplifies how a Z-order curve traverses a regular point grind in 3D space. By ordering the Gaussians according to their positions along a Z-order curve, spatial coherence can be exploited to improve the efficiency of run-length or predictive encoding techniques. This makes Z-order curves useful for spatial indexing and creating a coherent order for splats. In the context of sparse point clouds in 3DGS, Z-order curves are used to order splats based on their positions, which is memory efficient as it avoids allocating storage to empty regions. However, this approach has limitations in terms of maintaining neighborhood relationships: high-dimensional neighbors are not always close when mapped linearly, which affects the efficiency of operations that rely on true spatial proximity. More sophisticated methods, like Hilbert curves, can sometimes provide better locality for such datasets[7], offering improved efficiency in spatial indexing by ensuring that spatial neighbors are more likely to remain neighbors in the linear mapping.
Figure 6: Example of a regular 3D point grid traversed with Z-order curves.
3.2.5 Tri-planes and K-planes
Building on the idea of mapping multidimensional data into lower-dimensional grids, tri-plane factorization instead projects 3D positions onto three orthogonal 2D planes to store and retrieve attributes [6]. This creates a natural 2D organization, where Gaussians that are spatially close in 3D tend to occupy proximate coordinates on each plane, allowing 2D feature maps to be saved or edited like images while preserving some local neighborhood relationships. Rather than storing each primitive’s parameters individually, tri-plane factorization encodes attributes (e.g., color, opacity) as 2D feature images in the xy, xz, and yz planes. This process is visualized in Figure 7. The features are combined via simple operations such as concatenation, summation or multiplication and then decoded through a small MLP. Thus, the method is a hybrid of explicit storage (the planes) and implicit representation (the decoder).
Figure 7: The tri-plane model: A 3D query point is projected onto three orthogonal planes: xy, xz and yz. For 3DGS compression, the query can be the position of the primitive. The planes can store features, which are decoded into the other attributes (e.g. color, opacity) with a small MLP.
Compared to Z-order curves, which provide a single linear mapping of 3D points to 1D indices, tri-planes rely on three parallel 2D projections for more direct spatial correlation in each axis-aligned view. This makes it straightforward to compress or synthesize the plane images with standard 2D techniques. It also allows for synthesis of the feature planes with generative frameworks such as StyleGAN[20]. Furthermore, extension to higher-dimensional spaces yields K-Planes [15], where temporal or additional dimensions can be incorporated by factorizing them into extra 2D planes.
As with other dimensionality reduction methods, key parameters include the overall resolution of each plane (trading off granularity versus memory), how features from the three planes are aggregated, and the complexity of the MLP decoder that maps these aggregated features to the final per-Gaussian attributes. Similar to the Z-order approach, the aim is to exploit spatial coherence so that splats sharing similar properties lie close in the 2D feature maps. Because tri-plane factorization scales as $\mathcal{O}(N^{2})$ instead of $\mathcal{O}(N^{3})$, higher resolutions can be used than in dense voxel grids. This makes tri-plane factorization useful tool for 3DGS compression.
3.2.6 Self-Organizing Gaussians
@@
Another approach to map high-dimensional Gaussian parameters into 2D grids is through self-organization. This method was developed specifically for 3DGS compression[40]. Figure 8 depicts an example of the primitives of the Truck scene from the Tanks and Templesdataset reorganized into a 2D grid. The idea is based on the concept of Self-Organizing Maps[25], an unsupervised learning model that projects high-dimensional data onto a lower-dimensional grid, preserving the topological relationships between data points through competitive learning. Similar in motivation to Z-order curves and tri-planes, this 2D representation enables the use of standard image compression techniques, exploiting perceptual redundancies and ensuring local smoothness between neighboring splats. Unlike Z-order curves and tri-planes, which impose a fixed ordering, the Self-Organizing Gaussians representation is optimized for each scene, allowing neighbors in all dimensions to be modeled effectively. By organizing attributes into multiple data layers that share a consistent 2D layout, this technique provides a highly compressible and efficient representation of the original 3D scene.
Figure 8: The Gaussian primitives of the Truck scene mapped into a 2D layout using the Self-Organizing Gaussians approach[40]. An optimal arrangement is learned which preserves local relationships. This learned organization can utilize the available grid space more efficiently than fixed methods like Z-curves tri-planes.
3.2.7 Region Adaptive Hierarchical Transform (RAHT)
RAHT is a bottom-up transform that reorganizes color attributes along an octree (see also Section 3.2.1 on octrees). At each step, it pairs neighboring voxels according to their weights—the number of underlying voxels—and produces one low-frequency (DC) and one high-frequency coefficient[11]. The DC coefficients propagate upward for further merging, while the high-frequency coefficients are encoded immediately. Because the transform adapts to local density by adjusting the transform matrix to the weights, it effectively preserves small-scale details while reducing redundancy. Compared to graph-based transforms that require costly eigen-decompositions, RAHT relies on simple pairwise operations at each level, making it computationally efficient and suitable for real-time compression. After applying the transform, the resulting coefficients are quantized and entropy coded using an arithmetic coder with a Laplacian model, with sub-band parameters encoded with a Run-Length Golomb-Rice (RLGR) coder. This approach achieves compression performance comparable to state-of-the-art methods for point-cloud compression, but at significantly lower complexity. This hierarchical scheme is especially effective for attributes that exhibit strong local correlation, thus it can be adapted to encode attributes like color or density for 3D Gaussian splatting.
3.2.8 Discussion
The methods discussed form a toolkit to achieve compact and efficient 3D Gaussian Splatting, each with distinct advantages and disadvantages. Z-curves, tri-plane factorizations and 2D grid mappings reduce dimensional complexity in a simple and straightforward way. Fixed schemes like Z-order curves provide efficient ordering but may struggle with spatial coherence in sparse datasets. Adaptive techniques, such as anchor-based grouping or learned self-organizing mappings, can mitigate these issues by dynamically adjusting to scene-specific redundancies. Tri-planes may not use the available space efficiently, as empty space in 3D translates to empty patches in the planes. Self-Organizing Gaussians overcome this by learning a custom, non-linear mapping per scene. Hash-grids use adaptive, multiresolution indexing, which is fast and memory efficient, but is not an invertible mapping. Hash grids and tri-planes both store features, which need to be decoded into attributes with an MLP. While explicit methods allow direct access to Gaussian attributes, this implicit feature storage introduces compactness at the cost of requiring additional decoding.
Octrees adaptively allocate memory only to occupied regions, optimizing resource use for complex scenes, but require advanced attribute compression schemes like RAHT to exploit local correlations for compression. Anchor-based methods group Gaussians around representative points, reducing local redundancy efficiently, similar to vector quantization. While these methods are often considered independently, combining anchor-based representations with structured grids (e.g., hash grids, octrees) can enhance both local and global efficiency. Hybrid approaches allow for flexible spatial partitioning while maintaining fine-grained control over Gaussian placement.
All discussed methods are designed for efficient decoding, making them suitable for real-time applications. However, encoding complexity varies: Self-Organizing Gaussians require a computationally expensive optimization to determine element order, anchor-based methods involve clustering and optimization to define representative points, and feature-based methods like tri-planes and hash grids require a training process to establish features and MLP weights for decoding. This feature decoding adds a small computational overhead at decoding and rendering time.
Memory consumption at decoding and rendering time depends on the representation. For explicit methods, VRAM (GPU memory) usage is directly tied to the number of primitives, meaning that reducing their count through compaction (as discussed in the following sections) directly lowers memory requirements. In contrast, methods like hash grids, tri-planes, and vector quantization can use less memory overall by storing features in compact structures such as feature planes, hash tables, or codebooks. However, their memory usage is constant, meaning that reducing the number of primitives does not further decrease VRAM consumption. Depending on whether a system is memory- or compute-constrained, different trade-offs may be preferable. Large scenes often demand significant VRAM, making compact representations beneficial, while compute-constrained devices, such as VR headsets, may have sufficient memory but benefit from pre-decoding implicit representations into explicit attributes to reduce runtime computation.
@@ 3.4 Compaction
By assessing criteria such as gradients, pixel coverage, and saliency maps, ADC intelligently determines whether Gaussians should be cloned, split, or removed. This ensures that additional Gaussians are allocated where they are most needed, such as in high-frequency regions, while redundant or less impactful Gaussians are pruned. Figure 9 exemplifies the processes for (a) Gaussian primitive cloning, (b) Gaussian primitive splitting, and (c) Gaussian primitive pruning. Cloning produces an identical replica of the chosen Gaussian which is then incorporated into the scene. The splitting algorithm substitutes a Gaussian with a set number (default N = 2) of child Gaussians. These child Gaussians, after splitting, are situated within the bounds of the removed parent Gaussian. The children’s scaling is derived from the parent Gaussian’s scaling, reduced by a factor. The remaining attributes are directly inherited from the parent Gaussian.
Figure 9: From top to bottom, this figure shows the following processes: (a) Gaussian cloning, (b) Gaussian splitting, and (c) Gaussian pruning.
@@ 4 Efficient Strategies for Compression and Compaction in 3D Gaussian Splatting
As 3D Gaussian Splatting (3DGS) evolves as a prominent method for real-time scene rendering, its increasing adoption is challenged by substantial storage and computational requirements. This section addresses two essential optimization strategies: Compression (Sec. 4.1) and Compaction (Sec. 4.2). Compression reduces storage usage by employing methods such as vector quantization, which clusters similar Gaussian attributes to reduce redundancy, and structured representations that organize Gaussians into more compact forms like grids or anchor points. Compaction, on the other hand, focuses on optimizing the number and distribution of Gaussians, ensuring that only the most essential elements are retained while reducing unnecessary data. Figure 11 illustrates how Compression – attribute compression and structured representations – and Compaction intersect. Compaction here includes Gaussian pruning and densification approaches. Together, these strategies enhance the efficiency of 3DGS, making it more practical for a wide range of applications and devices.
@@ 4.1.1 Efficient Representation of Gaussian Attributes
Based on the assumption that many Gaussians share similar attributes, they can be quantized using Vector Quantization (VQ). Most commonly quantized attributes are then stored in codebooks based on K-means as in [42, 43, 44, 26, 12]. In LightGaussian [12], the authors employ VQ on Spherical Harmonics in combination with a significance score to omit VQ on Spherical Harmonics with higher significance. Similarly, [43] introduces a sensitivity-parameter which describes the sensitivity of the reconstruction quality to changes of the Gaussian attributes. This sensitivity measure is then used for sensitivity-aware VQ of Gaussian attributes (e.g.  Spherical Harmonics, shape). In [53], the authors categorize SH coefficients in degrees greater than 0 as unimportant and thus compressable through VQ. Furthermore for key attributes (opacity, scales, Euler angles, and 0-degree SH coefficients), Xie et al. reduce the entropy using RAHT before quantization. RDO-Gaussian [49] use entropy-constrained VQ with codebooks to quantize covariance and color parameters for a more compact representation. HAC [9] uses an anchor structure with associated Gaussians and introduces an Adaptive Quantization Module (AQM) designed to dynamically select quantization steps to facilitate entropy coding of the anchor attributes. The researchers in EAGLES [16] employ a methodology for quantizing attributes such as rotation, view-dependent color, and opacity by leveraging a latent vector associated with each attribute. This latent vector is integrated with a multilayer perceptron (MLP) decoder, which facilitates the decoding of latent representations into attribute values. To ensure differentiability during the training process, an additional latent approximation is preserved. A Straight-Through Estimator (STE) is then utilized to round the latent approximation before the propagation of gradients.
As described in Section 2, 3D Gaussians are characterized by 59 attributes, with the majority pertaining to color representation. Of these, the Spherical Harmonics (SH) coefficients encompass 48 attributes across three bands of SH. Notably, 45 out of the 48 SH coefficients are responsible for depicting view-dependent colors. To reduce storage, attribute pruning SH coefficients is a common strategy. In light of this, [44] introduces an adaptive adjustment mechanism for SH. The authors advocate for the calculation of an average transmittance per pixel and per view. In instances where there is minimal variation per view, this allows for the reduction of SH coefficients to lower bands or their complete removal, thereby significantly diminishing the memory footprint associated with each Gaussian primitive.